OpenStudy (anonymous):

Anyone to walk me through? The perimeter of a rectangle is 38 ft. The length is 3 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. Write the first equation. 2x+2y= Write the second equation. X=y+ What is the length? ( _) ft What is the width? ( ) ft

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the first equation is the perimeter...

OpenStudy (anonymous):



OpenStudy (anonymous):

2W+2L = perimeter

OpenStudy (anonymous):

the second equation asks about the relationship between L and W

OpenStudy (anonymous):

that's talked about in the problem...

OpenStudy (anonymous):

so the perimeter states it is 38ft (2w+2L)

OpenStudy (anonymous):

L= W+ ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Length= x, width=y, length longer than width means x=3+y perimeter = x+x+y+y , which is sum of all sides, using x=3+y in this equation perimeter = 6+y+y+y+y p=6+4y

OpenStudy (anonymous):

Ok so L=W+38? (Algebraic)?

OpenStudy (anonymous):

I am not paying attention to that person, I want to learn. If you're still willing to help me. :)

OpenStudy (anonymous):

well if you say length = x or L , then we need to understand length is 3 more than width which is say Y or W

OpenStudy (anonymous):

Confused.....

OpenStudy (anonymous):

please let me know where are you confused

OpenStudy (anonymous):

so L=W+38? got that far...

OpenStudy (anonymous):

I am sry, but L is not equal to w+38

OpenStudy (anonymous):

lets say Length = L and widht = w

OpenStudy (anonymous):

we know length = 3 more than width, right ? so L=3+W

OpenStudy (anonymous):

Perimeter is sum of all sides = length+length+width+width

OpenStudy (anonymous):

we have Perimeter =38 , 38=L+L+W+W, BUT WE KNOW that L=3+W , so put that in the former equation 38=3+W+3+W+W+W, we have 38= 6+4W

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

12.5? length and 9.5 for width. Is this correct?

OpenStudy (anonymous):

let meknow if there is reason for confusion

OpenStudy (anonymous):

I divided the perimeter by 4... 38/4 which gave me 9.5 then I added 3 to that which gave me 12.5 (because it 3 more than) Am I figuring this incorrectly?

OpenStudy (anonymous):

no no, we cant divide, we have to move 4 to left side of equation, take away 4 from 38

OpenStudy (anonymous):

I really have no idea what I am doing. Sigh.

OpenStudy (anonymous):

4 from 38 is 34

OpenStudy (anonymous):

I think everyone gave up on helping me. ....

OpenStudy (anonymous):

oops there was a typo, take away 6 from 38 and we get 32 32=4w, solve for w and w=8

OpenStudy (anonymous):

so w=8 l=11

OpenStudy (anonymous):

great

OpenStudy (anonymous):

thank you.